OpenStudy (anonymous):

2pi(integrate)(x(x^2-x^3)dx 2pi(integrate)(x(x^2-x^3)dx @Mathematics

OpenStudy (anonymous):

\[2\pi \int\limits_{}^{}x(x^2-x^3)dx\]

OpenStudy (anonymous):

\[2\pi * ( {x^4 \over 4} - {x^5 \over 5} + C )\]

OpenStudy (anonymous):

need steps? (not that there are many steps, but still) ;)

OpenStudy (anonymous):

yes please, but it is supposed to be \[(-\pi*x^4*(4x-5))/10\]

OpenStudy (anonymous):

or that is was what i got when i did it on the calculator

OpenStudy (anonymous):

forget the 2pi, since it's in front of the integral you can solve the integral and then add 2pi. You probably put the pi within the integral, the rest of the equation is the same....

OpenStudy (anonymous):

\[\int\limits_{}^{}{ x(x^2-x^3)} = \int\limits_{}^{}{x^3-x^4 } = \left[ {1\over 4}x^4 - {1\over5}x^5 + C \right]\]

OpenStudy (anonymous):

then multiply by 2pi

OpenStudy (anonymous):

ok. thanks :)